Justine is ice-skating at the Lloyd Center. What is her final velocity if she accelerates at a rate of


2.0 m/s2 for 3.5 seconds?


1
Expert's answer
2022-03-15T10:34:50-0400

The final velocity

"v_f=v_i+at\\
=0+2.0*3.5=7.0\:\rm m/s"